OpenStudy (anonymous):

Ralph is 3 times as old as Sara. In 4 years, Ralph will be only twice as old as Sara will be then. Find Ralph's age now. If x represents Sarah's age now, which of the following expressions represents Ralph's age in four years ? 3 x 6 x 2 x + 4 3 x + 4

OpenStudy (gorv):

ralph age =3x after 4 year sara's age=x+4 ralph's age=2*(x+4)=2x+8 which must be equal to =3x+4 2x+8=3x+4 8-4=3x-2x x=4

OpenStudy (gorv):

X+4 will represent ralph's age